Suppose that the number of bacteria in a colony triples every hour. Let Bn denote the number of bacteria in the colony after n hours.
(a) Set up a recurrence relation for Bn.
(b) If 100 bacteria are used to begin a new colony, how many bacteria will be in the colony in 10 hours?
1
Expert's answer
2020-10-13T14:48:25-0400

Solution to (a)

Let the starting value at hour 0 be B0

B1 =3B0

B2=3B1

B3 =3B2

Bn=3Bn-1

Thus the recurrence relation for Bn is

Bn=3Bn-1



Solution to (b)

B0= 100

"B_n = 3B_{n-1}"


"\therefore B_{10} = 3B_9 = 3(3B_8)=3(9B_7)=3(27B_6)""=3(81B_5)=3(243B_4)=3(729B_3)=3(2187B_2)""=3(6561B_1)=3(19683B_0)""\therefore B _{10} =59049B_0"

"=59049(100)=5904900"

Answer: 5904900 bacteria